I need help solving a math problem. I'm supposed to use the substitution method then elimination. I figured out the equations, but I'm confused on what to do next.
x + y + z = 34
1x + 5y + 20z = 237
y = z - 1

Answer by Boreal(15235)

x + y + z = 34
1x + 5y + 20z = 237
y = z - 1
Substitute y=z-1 into the first two.
x+z-1+z=34; x+2z=35
x+5z-5+20z=237; x+25z=242
===========================
x+2z=35
-x-25z=-242
-23z=-207
z=9
y=8
x=17
or (17,8,9)
=========================
check in 2
17+5(8)+20(9)=17+40+180=237